wiki:UC20180305

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 3/05/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Rachel Kuzio de Naray (GSU), Bill Ketzeback (APO), John Bally (CU), Chip Kobulnicky (UWy), Mary Beth Kaiser (JHU), Ben Williams (UW), Leslie Hebb (NAPG), Jon Holtzman (NMSU), Kal Kadlec (UW), Russet McMillan (APO), Anne Verbiscer (UVa), John Tobin (OU)


Identify minute-taker for this meeting: Nancy and Mark will do it

User feedback and comments from institutional representatives:

  • Colorado - nothing to report
  • Georgia State - nothing to report
  • JHU - nothing to report
  • NAPG - They have a user who is struggling with DIS data reduction given the scattered light. Russet recommended that they do sky subtraction far enough away from the target that the problem is minimized. Russet has started a conversation with Adam Kowalski at CU about this but hasn’t gotten far yet; Adam will get back to her before mid-March. NAPG also reported a concern about an ARCTIC filter configuration mixup (the obs-spec changed it to do focusing and the user forgot to change it back). Leslie requested that we ask the obs-specs to put settings back after focusing or remind the user to do so (the script to do this automatically doesn't work when the diffuser is installed). Russet reminded everyone of the importance of spot-checking one's data (although for this particular instance the seeing was so bad that this mistake might not have been noticed anyway).
  • NMSU - nothing to report
  • Oklahoma - nothing to report
  • Seattle - no report
  • UVa - input on time domain issue (discussed further below), nothing from users
  • Washington (given by Ben Williams) - nothing to report but this would have been Scott’s last telecon, UW is switching UC reps and the next person will be Eric Bellm
  • Wyoming - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 1/30/18 through 2/28/18

0) Overview

Reasonably good weather this period, a couple small storms and one that lasted about 4 days, providing the area with much needed rain. DIS blue camera vacuum is being serviced this week. The red camera recovered nicely from last month's servicing but blue only recovered somewhat so we decided to service it again.

1) Telescope

Telescope was operational with minimal issues. We did see some minor communications issues that we are monitoring and looking into some hardware issues. It appears to be time issues with TCC communications but we have not pinned it down. We know it takes a TCC restart to fix it.

2) Instruments

DIS: We pumped the red camera and blue camera last month and red scattered light is now quite good. Blue only improved by about 40% so it is on the pump this week in an effort to clear any contamination. DCam occasionally gives us issues. Original DCam is still at Andor for servicing. We have confirmation that it is now being worked on.

TripleSpec has been operational.

Agile has been operational this period.

Echelle is operational. Inter-Order Light is still improving slightly. Ecam shutter (new spare installed a couple weeks ago) is causing some issues. Original shutter has been refurbished and back in hand. There is also readout banding that is causing some issues. A replacement camera order from FingerLakes has been ordered, however FingerLakes lost the PO so delivery has been delayed.

NIC-FPS is operational.

ARCTIC is operational. We are working on improving filter wheel reliability. Users should verify that their desired filter is in place.

3) ARCSAT

M1 has been polished, M2 was refigured to correct the astigmatism. Both mirrors should go to get coated within the week. We have let a PO to have Ashdome look at dome servicing. Some testing on M2 focus seems to point to the algorithm DFM uses not being implemented at times or motor driver not responding.


Additional telescope and instrument discussion:

We've had some recent rain/snow at the site that is much needed, but we lost almost a week due to weather. Now we are getting into the spring windy season.

Bill pumped on the DIS blue camera last week to improve scattered light. That resulted in some improvement but not as much as in the red.

The echelle inter-order light remains above the thresholds and seems to be holding.

We are still in the process of getting Dcam repaired at Andor.

We are replacing Ecam but the purchase order got lost by the vendor so we probably won’t see it operational until the summer.

Regarding ARCSAT, the mirrors are nearly done. The astigmatism in M2 has been fixed although we haven’t seen any test data yet. We expect the mirrors to head off to get coated this week so the progress has been good. The engineering staff has had some issues with the secondary focus; they are talking with the DFM folks. They haven’t found any mechanical or electronic problems regarding focussing issues but the symptoms point towards their algorithm; we don’t know why it works sometimes and not others.

We noticed in January that the ARCTIC CCD temp was running cooler than was expected but don’t know why. This makes it more prone to dropping a quadrant, which then requires a daytime procedure in order to bring it back. This seems to be happening ~ once/month or less. We will see if the dewar comes back up in temperature as the weather starts to warm up.


Q1 3.5m scheduling:

There are 2 unassigned half nights remaining in the Q1 schedule. There was an expression of interest for one of them, and the other is 03/18B (dark).


Q2 3.5-m scheduling:

Lots of student groups want to come for site/training visits in Q2 (6 groups requested the same 2 weeks in June!). Russet is still finalizing the schedule and notes that some may not get their first choice of time.


Q2 ARCSAT scheduling:

There has been no call issued for Q2 proposals since it is unknown when the telescope will be back up and running. We will issue a call for Q2 if/when we know more.


TDAWG Update

The Time Domain Astronomy Working Group (TDAWG) report is finished and is posted on the 3.5m users' wiki. There is generally lots of support across ARC for doing time domain followup studies. The report made a series of recommendations, some of which would be fairly simple to implement and some of which are more complex and would require more careful planning. Two simple recommendations are 1) that the observatory should make time domain results more visible and available so that the broader ARC community is aware of what's going on and what came out of a particular target of opportunity observation, and 2) we should establish a repository where we can put calibration files for commonly used instrumental setups so that multiple investigators can share cals if necessary. There will be no more regularly scheduled TDAWG telecons, although they may meet occasionally to discuss some specific aspect of time domain work that the observatory is trying to implement. Anne provided feedback from UVa, which centered around questions of how time domain programs would be implemented in practice (e.g. how would users be informed, could there be situations where users could not be preempted under any circumstances, what would be the compensation for the observer who was interrupted, how is the time charged, how would someone who is being compensated be assured to get a night of comparable weather conditions to the night they gave up, etc.). Russet reviewed the current policies re: how time is requested, special protection policy, how time is compensated, etc. There is clearly more work to be done at the management level to flesh out some of the details of this. Nancy will send an email to 3.5m-general about the TDAWG report to notify all 3.5m users of its existence. She and Ben are still requesting feedback from users once they have had a chance to read the report. Nancy pointed out that not all recommendations in the report may be implemented - they are discussion points at the moment.


NA2 Baffle Project

John Wilson provided an email report, which Nancy relayed to the Users Committee. The current status is that the UVa team is currently comparing the Zemax and composite telescope SolidWorks models in order to identify discrepancies between the two for apertures and obscurations. Once they resolve the discrepancies and adopt a unified prescription they will next determine through Zemax the maximum field of view supported by the telescope without the NA2 baffle in place.


KOSMOS update

Nancy and Sarah discussed the current status of our efforts to acquire KOSMOS from KPNO. The ARC Board of Governors approved the plan and Sarah is now working with the KPNO folks on getting the equipment transferred. The KPNO staff are in the process of assembling all of the pieces (instrument plus all ancillary equipment, computers, etc.) for inventorying purposes so the contents of the long term loan are detailed explicitly. Jon mused whether in fact it was necessary to have the ARC Board of Governors approve the KOSMOS plan rather than the Users Committee since it's more of an operational plan. Nancy explained that since she originally thought that this agreement was going to be formalized through the MOU process (rather than a long-term loan) she needed the ARC Board's concurrence.


Training and Education Working Group

Nancy announced the establishment of the Training and Education Working Group, which will be a working group modeled after TDAWG whose purpose is to examine the use of the 3.5m telescope for student training and education. The charge of the TEWG is to provide an overview of institutional goals and practices, identify best practices and recommendations for student training that support different institutional goals, and provide APO policy recommendations (if needed) to optimize student use. Nancy announced the formation of this group today and will be soliciting participation from those faculty who have led student trips in the recent past as well as other stakeholders (APO staff, students, other interested ARC members).

The establishment of this working group is not in direct response to recent changes in operational structures at the site (e.g. reduced availability of housing and emergency services) that are necessitating the restriction of on-site group sizes. These recent issues have led to a current restriction (for Q2 2018) of a maximum group size of 12 people (10 students + 2 supervisors). The TEWG should discuss whether larger group sizes should be accommodated (i.e. whether they can still achieve the educational goals of the trips) and if so, other housing options should be explored.


ACTION ITEMS:

open from previous meetings:

  • All UC reps: coordinate with your institutional scheduler to get a sense from your users concerning the degree to which they would be willing to be preempted for time domain followup work, and/or whether they would be interested in pooling time across ARC institutions to dedicate to time domain followup observations. Status: CLOSED. Ben has now heard from everyone, including UVa on today's call.

new from this meeting:

  • Nancy will notify all 3.5m users (through an email to 35m-general) about the availability of the TDAWG report.
  • Nancy will begin recruiting members for the TEWG.

Next meeting:

The meeting schedule for the spring semester will be the first Monday of the month at 11:00 am Mountain Time. The next call will thus be April 2, 2018.

Future meetings are scheduled for:

  • May 7
  • June 4
  • July 9

We will evaluate everyone's availability (as well as the telecon schedule) once we get into summer months since we will need to contend with summer travel schedules as well as summer shutdown activities.

Note: See TracWiki for help on using the wiki.